# The single edge notch fracture test for viscoelastic elastomers

### v1

@@ 4.1 Constant viscosity
In addition to the results for the critical global stretch $\mathrm{\Lambda}_{c}$ and the critical global stress $S_{c}$, it is instructive to plot the results for the corresponding critical tearing energy (21). Figure 7 includes such a plot in terms of the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$ for the three material constants $\beta_{2}=1,2.5,7.5$. The results show the typical “S” shape found in other types of fracture tests, such as the trousers fracture testGreensmith and Thomas (1955), with the caveat that they are not strictly monotonic. Specifically, as $\dot{\mathrm{\Lambda}}_{0}\searrow 0$, for sufficiently slow global stretch rates, $T_{c}\searrow G_{c}$. As $\dot{\mathrm{\Lambda}}_{0}$ increases, so does $T_{c}$, although not necessarily monotonically. As $\dot{\mathrm{\Lambda}}_{0}\nearrow+\infty$, for sufficiently fast global stretch rates, $T_{c}$ approaches an asymptotic value, $T_{c}^{\infty}$ say. The transition of $T_{c}$ from its minimum value $G_{c}$ to its asymptotic value $T_{c}^{\infty}$ is controlled by both the non-equilibrium elasticity of the elastomer and its viscosity. Given that $G_{c}=150$ N/m and $T_{c}^{\infty}>3000$ N/m, the results in Fig. 7 make it plain that the critical tearing energy $T_{c}$ is primarily a direct manifestation of the viscoelastic behavior of the elastomer, and not of its fracture behavior. In other words, in view of the representation (3), the value of $T_{c}$ is mostly a measure of how much energy is dissipated by viscous deformation and not by the creation of surface.
@@ 4.2 Shear-thinning viscosity
Figure 8 presents results analogous to those presented in Fig. 7 for elastomers whose viscosity is not constant but instead decreases with deformation rate, that is, a viscosity that exhibits shear thinning; the corresponding force-deformation results analogous to those in Fig. 6 are presented in Fig. S2 of the Supplementary Material. Precisely, the results pertain to specimens with initial crack size $A=2$ mm, elastomers with non-equilibrium elasticity of weaker, equal, and stronger growth conditions (23), with $\nu=\mu^{{\rm NEq}}=20\mu^{{\rm Eq}}=2$ MPa, than their equilibrium elasticity, and shear-thinning viscosity (25)2 with $\eta_{0}=\nu\tau_{0}=2$ MPa s. Much like the results in Fig. 7, the results in Fig. 8 correspond to the representative case when the critical energy release rate is $G_{c}=150$ N/m.
@@ 4.3 Deformation-dependent viscosity
Figure 9 presents results analogous to those presented in Figs. 7 and 8 for elastomers whose viscosity increases with increasing deformation; the corresponding force-deformation results analogous to those in Fig. 6 are presented in Fig. S3 of the Supplementary Material. In particular, the results pertain to specimens with initial crack size $A=2$ mm, elastomers with non-equilibrium elasticity of weaker, equal, and stronger growth conditions (23), with $\nu=\mu^{{\rm NEq}}=20\mu^{{\rm Eq}}=2$ MPa, than their equilibrium elasticity, and deformation-dependent viscosity (25)3 with $\eta_{0}=\nu\tau_{0}=2$ MPa s. Again, much like the results in the two preceding subsections, the results in Fig. 9 correspond to the representative case when the critical energy release rate is $G_{c}=150$ N/m.
@@
The classical formula (4) of Rivlin and ThomasRivlin and Thomas (1953) — together with the approximation (5) for the numerical factor $K(\mathrm{\Lambda})$ that emerged from the works of GreensmithGreensmith (1963) and LindleyLindley (1972) — is commonly used well beyond its original scope (see Subsection 1.1 above) in the experimental literature so as to estimate the values of the critical tearing energy $T_{c}$ from single edge notch fracture tests. In the sequel, we show how such an overextended use can lead to significant errors in the estimation of $T_{c}$. We do so by confronting the formula (4) directly with an exact solution from our 3D full-field simulations for viscoelastic elastomers.
Remark 3
@@
Figure 10 provides a comparison between the energy release rate $-\partial\mathcal{W}/\partial\mathrm{\Gamma}_{0}$ generated by the classical formula (4), with the numerical factor (5), and the exact result (21) determined from 3D full-field simulations of a single edge notch test, carried out at the global stretch rate $\dot{\mathrm{\Lambda}}_{0}=10^{0}$ s-1, for a specimen with pre-existing crack of length $A=2$ mm and an elastomer with shear-thinning viscosity (25)2, $\eta_{0}=\nu\tau_{0}=2$ MPa s, and non-equilibrium elasticity of equal growth conditions (23)2, with $\nu=\mu^{{\rm NEq}}=20\mu^{{\rm Eq}}=2$ MPa, than its equilibrium elasticity.
The comparison in Fig. 10 makes it clear that the classical formula (4) of Rivlin and ThomasRivlin and Thomas (1953) can be greatly inaccurate, increasingly more so for larger global stretches $\mathrm{\Lambda}$, when used for elastomers that dissipate energy by deformation and for cracks that are not much smaller than the length and the width of the specimen.
@@ 5 Comparisons with experiments on an acrylate elastomer
As alluded to in the Introduction, since the pioneering work of BusseBusse (1934) in the 1930s, rather remarkably, only a handful of the numerous single edge notch fracture tests that have been reported in the literature have been carried out at different loading rates $\dot{\mathrm{\Lambda}}_{0}$. In this section, we deploy the Griffith criticality condition (2) to explain one of them, those of Le MennMenn (2022) on an acrylate elastomer.
As part of their doctoral thesis, Le MennMenn (2022) carried out a series of single edge notch tests on dog-bone specimens of gauge length $L=15$ mm, width $H=4$ mm, thickness $B\in[0.2,0.8]$ mm, and crack size $A=1$ mm made of acrylate elastomers (for use as coatings of optical fibers) at room temperature and four different constant global stretch rates, approximately given by $\dot{\mathrm{\Lambda}}_{0}=6.7\times 10^{-5},6.7\times 10^{-4},6.7\times 10^{-3},3.3\times 10^{-2}$ s-1; see Appendix B.4 in their work. Out of all the various compositions that Le MennMenn (2022) considered, we focus here on the one that she labeled OL HEA 30% PEA for which she provided experimental data for the elastic response of the elastomer in the form of stress-stretch data for uniaxial and equi-biaxial tension carried out at a constant stretch rate of about $\dot{\lambda}_{0}=6.7\times 10^{-3}$ s-1, sufficiently slow that viscous effects were not observed; see Chapter 3 in their work.
Figure 11: Comparison between the stress-stretch response described by the viscoelastic model (11)-(12), with the material constants in Table 1, and the experimental data reported by Le MennMenn (2022) for the acrylate elastomer OL HEA 30% PEA subjected to uniaxial and equi-biaxial tension applied at a slow constant stretch rate $\dot{\lambda}_{0}$ at room temperature.
Figure 12: (a) Comparison between the force-deformation response predicted theoretically and the experimental data reported by Le MennMenn (2022) for single edge notch fracture tests on the acrylate elastomer OL HEA 30% PEA carried out at four different global stretch rates $\dot{\mathrm{\Lambda}}_{0}$. (b, c) Corresponding comparison between the critical global stretch $\mathrm{\Lambda}_{c}$ and the critical global stress $S_{c}$ at which fracture nucleates predicted by the Griffith criticality condition (2) and the experimental results.
@@
The fitting of the material constants $\mu_{1}$, $\mu_{2}$, $\alpha_{1}$, $\alpha_{2}$, which, again, describe the equilibrium elasticity in the viscoelastic model (11)-(12), to the uniaxial and equi-biaxial data reported by Le MennMenn (2022) yields the values listed in the first column of Table 1. As seen from the comparisons presented in Fig. 11, the constitutive relation (11)-(12) with such material constants describes reasonably well the experimentally measured response at states of thermodynamic equilibrium.
@@
[…]
$\alpha_{1}=4$ $\beta_{1}=-0.5$ $\eta_{\infty}=0$
[…]
@@
It remains to determine the materials constants $\nu_{1}$, $\nu_{2}$, $\beta_{1}$, $\beta_{2}$ for the non-equilibrium elasticity and the viscosity function $\eta({\bf C},{\bf C}^{v})$. Given the lack of direct experimental data beyond that presented in Fig. 11 for the viscoelastic behavior of OL HEA 30% PEA, we turn to remarking that the force-deformation response of the specimens in the single edge notch fracture tests reported by Le MennMenn (2022), reproduced in Fig. 12(a), differ only on the value of the critical global stretch $\mathrm{\Lambda}_{c}$ — and hence also on the value of the critical global stress $S_{c}$ — at which fracture nucleates. In other words, the force-deformation response of the specimens appears to be independent of the loading rate $\dot{\mathrm{\Lambda}}_{0}$, except for the fact that $\mathrm{\Lambda}_{c}$ and $S_{c}$ increase with increasing $\dot{\mathrm{\Lambda}}_{0}$. Given that the stretch and stretch rates around the crack front in single edge notch fracture tests are locally very large, such a behavior may be attributed to a viscosity that is strongly nonlinear at large stretches and stretch rates, but one that is negligible otherwise. With this mechanism in mind, we consider a viscosity function of the piecewise form
$\eta({\bf C},{\bf C}^{v})=\left\{\begin{array}[]{ll}\eta_{0},&\,\textbf{X}\in\mathrm{\Omega}_{0}^{\rm I}\vspace{0.2cm}\\ \eta_{\infty}+\dfrac{\eta_{0}-\eta_{\infty}}{1+\left(K_{2}\mathcal{J}_{2}^{\rm{NEq}}\right)^{\gamma_{2}}},&\,\textbf{X}\in\mathrm{\Omega}_{0}^{\rm II}\vspace{0.2cm}\\ \eta_{\infty},&\,\textbf{X}\in\mathrm{\Omega}_{0}\setminus(\mathrm{\Omega}_{0}^{\rm I}\cup\mathrm{\Omega}_{0}^{\rm II})\end{array}\right.$ (27)
@@
Given the viscosity function (27), a search of the material constants $\nu_{1}$, $\nu_{2}$, $\beta_{1}$, $\beta_{2}$, $\eta_{0}$, $\eta_{\infty}$, $K_{2}$, $\gamma_{2}$ in the viscoelastic model (11)-(12) that are consistent with the data of Le MennMenn (2022) yields the values listed in the second and third columns of Table 1.
Figure 13: Schematics of the small regions $\mathrm{\Omega}_{0}^{\rm{I}}$ and $\mathrm{\Omega}_{0}^{\rm{II}}$ around the crack front in the single edge notch fracture tests carried out by Le MennMenn (2022) wherein the viscoelastic behavior of the acrylate elastomer OL HEA 30% PEA is described by the nonlinear viscosity (27)1,2.
@@
which is consistent with the estimate provided by Le MennMenn (2022); see Table 3.7 in their work.
At this stage, having calibrated the viscoelastic behavior and the critical energy release rate of the elastomer, we are ready to deploy the Griffith criticality condition (2) to confront its predictions with the single edge notch fracture tests reported by Le MennMenn (2022). Figures 12(b,c) do so for the critical global stretch $\mathrm{\Lambda}_{c}$ and the critical global stress $S_{c}$ at which fracture nucleates as functions of the applied global stretch rate $\dot{\mathrm{\Lambda}}_{0}$. The error bars for the experimental data correspond to the neighborhood over which the force-deformation response of the specimens approaches its peak value and hence over which the onset of fracture is bound to occur.
@@ 6 Summary and final comments
Additionally, the parametric study presented in Section 4 has served to establish that the classical formula (4) of Rivlin and ThomasRivlin and Thomas (1953), together with the approximation (5) for the numerical factor $K(\mathrm{\Lambda})$ due to GreensmithGreensmith (1963) and LindleyLindley (1972), can be greatly inaccurate when used to estimate the critical tearing energy $T_{c}$ for elastomers that dissipate energy by deformation and for cracks that are not much smaller than the length and the width of the specimen. This result should not come as a surprise, since Rivlin and ThomasRivlin and Thomas (1953) derived the formula (4) precisely under the premise that energy dissipation (viscous or otherwise) by deformation is absent and that the pre-existing crack is much smaller than the length and the width of the specimen, but much larger than its thickness.
The comparisons with experiments presented in Section 5 have also served to add to previous validation results for other tests and materialsShrimali and Lopez-Pamies (2023a, b, c) and thus they have provided further evidence that the Griffith criticality condition (2) may indeed be the condition that governs when large cracks grow in viscoelastic elastomers subjected to arbitrary quasi-static loading conditions.
To conclusively establish whether the Griffith criticality condition (2) is “the law of the land”, more comparisons with experiments on different elastomers under different types of loadings would be essential. In this regard, we emphasize that a key implication of (2) is that any rate effects on the nucleation of fracture from large pre-existing cracks in elastomers are due to their nonlinear viscoelastic behavior. This, we hope, will motivate experimentalists and theoreticians alike to thoroughly characterize the nonlinear viscoelasticity of elastomers over large ranges of stretches and stretch rates — and not just within the classical limited setting of small deformations, as most often donePersson et al. (2024), which is insufficient to analyze fracture — so as to help establishing whether all rate effects observed in the fracture of elastomers are indeed of a viscous nature.

### v4

@@ 4.1 Constant viscosity
In addition to the results for the critical global stretch $\mathrm{\Lambda}_{c}$ and the critical global stress $S_{c}$, it is instructive to plot the results for the corresponding critical tearing energy (42). Figure 7 includes such a plot in terms of the global stretch rate $\dot{\mathrm{\Lambda}}_{0}$ for the three material constants $\beta_{2}=1,2.5,7.5$. The results show the typical “S” shape found in other types of fracture tests, such as the trousers fracture test [Greensmith55], with the caveat that they are not strictly monotonic. Specifically, as $\dot{\mathrm{\Lambda}}_{0}\searrow 0$, for sufficiently slow global stretch rates, $T_{c}\searrow G_{c}$. As $\dot{\mathrm{\Lambda}}_{0}$ increases, so does $T_{c}$, although not necessarily monotonically. As $\dot{\mathrm{\Lambda}}_{0}\nearrow+\infty$, for sufficiently fast global stretch rates, $T_{c}$ approaches an asymptotic value, $T_{c}^{\infty}$ say. The transition of $T_{c}$ from its minimum value $G_{c}$ to its asymptotic value $T_{c}^{\infty}$ is controlled by both the non-equilibrium elasticity of the elastomer and its viscosity. Given that $G_{c}=150$ N/m and $T_{c}^{\infty}>3000$ N/m, the results in Fig. 7 make it plain that the critical tearing energy $T_{c}$ is primarily a direct manifestation of the viscoelastic behavior of the elastomer, and not of its fracture behavior. In other words, in view of the representation (5), the value of $T_{c}$ is mostly a measure of how much energy is dissipated by viscous deformation and not by the creation of surface.
@@ 4.2 Shear-thinning viscosity
Figure 8 presents results analogous to those presented in Fig. 7 for elastomers whose viscosity is not constant but instead decreases with deformation rate, that is, a viscosity that exhibits shear thinning; the corresponding force-deformation results analogous to those in Fig. 6 are presented in Fig. 15 in the Appendix. Precisely, the results pertain to specimens with initial crack size $A=2$ mm, elastomers with non-equilibrium elasticity of weaker, equal, and stronger growth conditions (46), with $\nu=\mu^{{\rm NEq}}=20\mu^{{\rm Eq}}=2$ MPa, than their equilibrium elasticity, and shear-thinning viscosity (48)2 with $\eta_{0}=\nu\tau_{0}=2$ MPa s. Much like the results in Fig. 7, the results in Fig. 8 correspond to the representative case when the critical energy release rate is $G_{c}=150$ N/m.
@@ 4.3 Deformation-dependent viscosity
Figure 9 presents results analogous to those presented in Figs. 7 and 8 for elastomers whose viscosity increases with increasing deformation; the corresponding force-deformation results analogous to those in Fig. 6 are presented in Fig. 16 in the Appendix. In particular, the results pertain to specimens with initial crack size $A=2$ mm, elastomers with non-equilibrium elasticity of weaker, equal, and stronger growth conditions (46), with $\nu=\mu^{{\rm NEq}}=20\mu^{{\rm Eq}}=2$ MPa, than their equilibrium elasticity, and deformation-dependent viscosity (48)3 with $\eta_{0}=\nu\tau_{0}=2$ MPa s. Again, much like the results in the two preceding subsections, the results in Fig. 9 correspond to the representative case when the critical energy release rate is $G_{c}=150$ N/m.
@@
The classical formula (6) of [RT53] — together with the approximation (7) for the numerical factor $K(\mathrm{\Lambda})$ that emerged from the works of [Greensmith63] and [Lindley72] — is commonly used well beyond its original scope (see Subsection 1.1 above) in the experimental literature so as to estimate the values of the critical tearing energy $T_{c}$ from single edge notch fracture tests. In the sequel, we show how such an overextended use can lead to significant errors in the estimation of $T_{c}$. We do so by confronting the formula (6) directly with an exact solution from our 3D full-field simulations for viscoelastic elastomers.
Remark 4.1.
@@
Figure 10 provides a comparison between the energy release rate $\partial\mathcal{W}/\partial\mathrm{\Gamma}_{0}$ generated by the classical formula (6), with the numerical factor (7), and the exact result (42) determined from 3D full-field simulations of a single edge notch test, carried out at the global stretch rate $\dot{\mathrm{\Lambda}}_{0}=10^{0}$ s-1, for a specimen with pre-existing crack of length $A=2$ mm and an elastomer with shear-thinning viscosity (48)2, $\eta_{0}=\nu\tau_{0}=2$ MPa s, and non-equilibrium elasticity of equal growth conditions (46)2, with $\nu=\mu^{{\rm NEq}}=20\mu^{{\rm Eq}}=2$ MPa, than its equilibrium elasticity.
The comparison in Fig. 10 makes it clear that the classical formula (6) of [RT53] can be greatly inaccurate, increasingly more so for larger global stretches $\mathrm{\Lambda}$, when used for elastomers that dissipate energy by deformation and for cracks that are not much smaller than the length and the width of the specimen.
@@ 5 Comparisons with experiments on an acrylate elastomer
As alluded to in the Introduction, since the pioneering work of [Busse34] in the 1930s, rather remarkably, only a handful of the numerous single edge notch fracture tests that have been reported in the literature have been carried out at different loading rates $\dot{\mathrm{\Lambda}}_{0}$. In this section, we deploy the Griffith criticality condition (2) to explain one of them, those of [LeMenn22] on an acrylate elastomer.
As part of their doctoral thesis, [LeMenn22] carried out a series of single edge notch tests on dog-bone specimens of gauge length $L=15$ mm, width $H=4$ mm, thickness $B\in[0.2,0.8]$ mm, and crack size $A=1$ mm made of acrylate elastomers (for use as coatings of optical fibers) at room temperature and four different constant global stretch rates, approximately given by $\dot{\mathrm{\Lambda}}_{0}=6.7\times 10^{-5},6.7\times 10^{-4},6.7\times 10^{-3},3.3\times 10^{-2}$ s-1; see Appendix B.4 in their work. Out of all the various compositions that [LeMenn22] considered, we focus here on the one that she labeled OL HEA 30% PEA for which she provided experimental data for the elastic response of the elastomer in the form of stress-stretch data for uniaxial and equi-biaxial tension carried out at a constant stretch rate of about $\dot{\lambda}_{0}=6.7\times 10^{-3}$ s-1, sufficiently slow that viscous effects were not observed; see Chapter 3 in their work.
Figure 11: Comparison between the stress-stretch response described by the viscoelastic model (24)-(25), with the material constants in Table 1, and the experimental data reported by [LeMenn22] for the acrylate elastomer OL HEA 30% PEA subjected to uniaxial and equi-biaxial tension applied at a slow constant stretch rate $\dot{\lambda}_{0}$ at room temperature.
Figure 12: (a) Comparison between the force-deformation response predicted theoretically and the experimental data reported by [LeMenn22] for single edge notch fracture tests on the acrylate elastomer OL HEA 30% PEA carried out at four different global stretch rates $\dot{\mathrm{\Lambda}}_{0}$. (b, c) Corresponding comparison between the critical global stretch $\mathrm{\Lambda}_{c}$ and the critical global stress $S_{c}$ at which fracture nucleates predicted by the Griffith criticality condition (2) and the experimental results.
@@
The fitting of the material constants $\mu_{1}$, $\mu_{2}$, $\alpha_{1}$, $\alpha_{2}$, which, again, describe the equilibrium elasticity in the viscoelastic model (24)-(25), to the uniaxial and equi-biaxial data reported by [LeMenn22] yields the values listed in the first column of Table 1. As seen from the comparisons presented in Fig. 11, the constitutive relation (24)-(25) with such material constants describes reasonably well the experimentally measured response at states of thermodynamic equilibrium.
@@
$\mu_{1}=$0.400\text{\,}\mathrm{MPa}$\$ $\nu_{1}=$5.19\text{\,}\mathrm{MPa}$\$ $\eta_{0}=$80\text{\,}\mathrm{MPa}\text{\,}\mathrm{s}$\$
$\alpha_{1}=0.0488$ $\beta_{1}=-0.5$ $\eta_{\infty}=0$
$\mu_{2}=$0.119\text{\,}\mathrm{MPa}$\$ $\nu_{2}=0$ $K_{2}=$0.01\text{\,}\mathrm{MPa}$^{-2}$
@@
It remains to determine the materials constants $\nu_{1}$, $\nu_{2}$, $\beta_{1}$, $\beta_{2}$ for the non-equilibrium elasticity and the viscosity function $\eta({\bf C},{\bf C}^{v})$. Given the lack of direct experimental data beyond that presented in Fig. 11 for the viscoelastic behavior of OL HEA 30% PEA, we turn to remarking that the force-deformation response of the specimens in the single edge notch fracture tests reported by [LeMenn22], reproduced in Fig. 12(a), differ only on the value of the critical global stretch $\mathrm{\Lambda}_{c}$ — and hence also on the value of the critical global stress $S_{c}$ — at which fracture nucleates. In other words, the force-deformation response of the specimens appears to be independent of the loading rate $\dot{\mathrm{\Lambda}}_{0}$, except for the fact that $\mathrm{\Lambda}_{c}$ and $S_{c}$ increase with increasing $\dot{\mathrm{\Lambda}}_{0}$. Given that the stretch and stretch rates around the crack front in single edge notch fracture tests are locally very large, such a behavior may be attributed to several plausible mechanisms. One possibility is that it is a direct manifestation of the strength surface of the elastomer; see, e.g., Sections 2.2.1 and 5.1 in [KBLP25]. Another possibility is that the viscosity of the elastomer is one that is strongly nonlinear at large stretches and stretch rates, but negligible otherwise. In this work, we explore this latter mechanism. Accordingly, we consider a viscosity function of the piecewise form
$\eta({\bf C},{\bf C}^{v})=\left\{\hskip-1.42271pt\begin{array}[]{ll}\eta_{0},&\,\textbf{X}\in\mathrm{\Omega}_{0}^{\rm I}\vskip 5.69046pt\\ \eta_{\infty}+\dfrac{\eta_{0}-\eta_{\infty}}{1+\left(K_{2}\mathcal{J}_{2}^{\rm{NEq}}\right)^{\gamma_{2}}},&\,\textbf{X}\in\mathrm{\Omega}_{0}^{\rm II}\vskip 5.69046pt\\ \eta_{\infty},&\,\textbf{X}\in\mathrm{\Omega}_{0}\hskip-0.85355pt\setminus\hskip-0.56917pt(\mathrm{\Omega}_{0}^{\rm I}\cup\mathrm{\Omega}_{0}^{\rm II})\end{array}\right$. (51)
@@
Given the viscosity function (51), a search of the material constants $\nu_{1}$, $\nu_{2}$, $\beta_{1}$, $\beta_{2}$, $\eta_{0}$, $\eta_{\infty}$, $K_{2}$, $\gamma_{2}$ in the viscoelastic model (24)-(25) that are consistent with the data of [LeMenn22] yields the values listed in the second and third columns of Table 1.
Figure 13: Schematics of the small regions $\mathrm{\Omega}_{0}^{\rm{I}}$ and $\mathrm{\Omega}_{0}^{\rm{II}}$ around the crack front in the single edge notch fracture tests carried out by [LeMenn22] wherein the viscoelastic behavior of the acrylate elastomer OL HEA 30% PEA is described by the nonlinear viscosity (51)1,2.
@@
which is consistent with the estimate provided by [LeMenn22]; see Table 3.7 in their work.
At this stage, having calibrated the viscoelastic behavior and the critical energy release rate of the elastomer, we are ready to deploy the Griffith criticality condition (2) to confront its predictions with the single edge notch fracture tests reported by [LeMenn22]. Figures 12(b,c) do so for the critical global stretch $\mathrm{\Lambda}_{c}$ and the critical global stress $S_{c}$ at which fracture nucleates as functions of the applied global stretch rate $\dot{\mathrm{\Lambda}}_{0}$. The error bars for the experimental data correspond to the neighborhood over which the force-deformation response of the specimens approaches its peak value and hence over which the onset of fracture is bound to occur.
@@ 6 Summary and final comments
Additionally, the parametric study presented in Section 4 has served to establish that the classical formula (6) of [RT53], together with the approximation (7) for the numerical factor $K(\mathrm{\Lambda})$ due to [Greensmith63] and [Lindley72], can be greatly inaccurate when used to estimate the critical tearing energy $T_{c}$ for elastomers that dissipate energy by deformation and for cracks that are not much smaller than the length and the width of the specimen. This result should not come as a surprise, since [RT53] derived the formula (6) precisely under the premise that energy dissipation (viscous or otherwise) by deformation is absent and that the pre-existing crack is much smaller than the length and the width of the specimen, but much larger than its thickness.
The comparisons with experiments presented in Section 5 have also served to add to previous validation results for other tests and materials [SLP23a, SLP23b, SLP23c] and thus they have provided further evidence that the Griffith criticality condition (2) may indeed be the condition that governs when large cracks grow in viscoelastic elastomers subjected to arbitrary quasi-static loading conditions.
To conclusively establish whether the Griffith criticality condition (2) is “the law of the land”, more comparisons with experiments on different elastomers under different types of loadings would be essential. In this regard, we emphasize that a key implication of (2) is that any rate effects on the nucleation of fracture from large pre-existing cracks in elastomers are due to their nonlinear viscoelastic behavior. This, we hope, will motivate experimentalists and theoreticians alike to thoroughly characterize the nonlinear viscoelasticity of elastomers over large ranges of stretches and stretch rates — and not just within the classical limited setting of small deformations, as most often done [Persson2024], which is insufficient to analyze fracture — so as to help establishing whether all rate effects observed in the fracture of elastomers are indeed of a viscous nature.